OpenStudy (anonymous):

What is the mean or average of the five measurements below? {0.5 cm, 0.5 cm, 0.8 cm, 0.7 cm, 0.7 cm} 1.6 cm 0.5 cm 0.64 cm 3.2 cm

OpenStudy (homeworkhelper099):

Add and Divide?

OpenStudy (anonymous):

Add and Divide what?

OpenStudy (homeworkhelper099):

Add them all up them divide.

OpenStudy (homeworkhelper099):

But the answer is : 0.64 cm

OpenStudy (anonymous):

You're correct

OpenStudy (homeworkhelper099):

:)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

What is the mean of the following set of numbers? 8.025 9.3 3.8 12.035

OpenStudy (anonymous):

What is the mean of the following set of numbers? 8.025 9.3 3.8 12.035

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?3 \,\frac{3}{8}\,,\;9.3,\;11.4

OpenStudy (homeworkhelper099):

Is this just like your last one?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Thats not one of the answers.

OpenStudy (homeworkhelper099):

13.85?

OpenStudy (anonymous):

A. 8.025 B. 9.3 C. 3.8 D. 12.035

OpenStudy (homeworkhelper099):

what's the question???

OpenStudy (anonymous):

What is the mean of the following set of numbers? https://media.studyisland.com/cgi-bin/mimetex.cgi?3 \,\frac{3}{8}\,,\;9.3,\;11.4

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?3 \,\frac{3}{8}\,,\;9.3,\;11.4

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Let me copy it real quik

OpenStudy (homeworkhelper099):

ok

OpenStudy (anonymous):

there

OpenStudy (homeworkhelper099):

sorry buddy, i don't know.

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

I think it's 8.025

OpenStudy (anonymous):

Yeah it is